OpenStudy (anonymous):

it is possible to have both Summation a_n and Summation b_n be divergent series and yet have Summation from n=1 to inf (a_n +- b_n) be a convergent series. What does this statement means? Does it mean that if both Summation a_n and Summation b_n are divergent series then is it possible to Summation from n=1 to inf (a_n +- b_n) to be a convergence?

OpenStudy (zarkon):

if \(a_n=1\) and \(b_n=-1\) \[\sum_{n=1}^{\infty}a_n\text{ diverges}\] and \[\sum_{n=1}^{\infty}b_n\text{ diverges}\] but \[\sum_{n=1}^{\infty}(a_n+b_n)=0\]

OpenStudy (anonymous):

OK